The length of the longer leg of a right triangle is
13cm
more than three times the length of the shorter leg. The length of the hypotenuse is
14cm
more than three times the length of the shorter leg. Find the side lengths of the triangle.

Answer by ewatrrr(24785)

.
(3x + 13)^2 + x^2 = (3x+14)^2
3x^2 + 78x +169 + x^2 = 3x^2 + 84x + 196
x^2 - 6x - 27 = 0
(x+3)(x-9) = 0
x = 9
hypotenuse = 41
....checking...
40^2 + 9^2 = 41^2